OpenStudy (anonymous):

If sin A=3/5, find sin 2A and sin 4A

OpenStudy (shayaan_mustafa):

Hi lilfayfay! :) There are two ways to find these values. One by the formula of double angle and other is simple without formula. Which way do yo prefer buddy?

OpenStudy (anonymous):

\[A=\sin ^{-1}\left(\frac{3}{5}\right) \]\[\{\sin (2 A),\sin (4 A)\}=\left\{\sin \left(2 \sin ^{-1}\left(\frac{3}{5}\right)\right),\sin \left(4 \sin ^{-1}\left(\frac{3}{5}\right)\right)\right\}=\{0.96,0.5376\} \]